### v3

# Enhancing Few-Shot Vision-Language Classification with Large Multimodal Model Features

Abstract
Generative Large Multimodal Models (LMMs) like LLaVA and Qwen-VL excel at a wide variety of vision-language (VL) tasks. Despite strong performance, LMMs’ generative outputs are not specialized for vision-language classification tasks (i.e., tasks with vision-language inputs and discrete labels) such as image classification and multiple-choice VQA. One key challenge in utilizing LMMs for these tasks is the extraction of useful features from generative LMMs. To overcome this, we propose an approach that leverages multimodal feature extraction from the LMM’s latent space. Toward this end, we present Sparse Attention Vectors (SAVs)—a finetuning-free method that leverages sparse attention head activations (fewer than 5% of the heads) in LMMs as strong feature representations. With only few-shot examples, SAVs demonstrate state-of-the-art performance compared to a variety of few-shot and finetuned baselines on a collection of vision-language classification tasks. Our experiments also imply that SAVs can scale in performance with additional examples and generalize to similar tasks, establishing SAVs as both effective and robust multimodal feature representations.
1 Introduction
Generative Large Multimodal Models (LMMs) such as GPT-4V [72], LLaVA [59, 58], and QwenVL [3] demonstrate state-of-the-art performance on open-ended vision-language (VL) tasks like image captioning [53, 105], visual question answering [2, 31, 42], and language grounding [37, 64]. However, despite their remarkable performance on generative tasks, these models struggle on vision-language classification tasks, where responses are a discrete set of labels [109, 7]. Indeed, LMMs with billions of parameters and trained on trillions more tokens of data underperform smaller encoder VLMs like CLIP and SigLIP [109, 7] and even classical machine learning methods [5] on image classification tasks. One reason that CLIP-like models excel at classification is that is straightforward to extract visual (or textual) features from visual (or textual) encoders. However, such models are not equipped to process joint vision-language inputs, as LMMs can. Our goal is to extract CLIP-like features from generative LMMs, enabling downstream discrete labeling tasks such as classification and multiple-choice VQA.
Figure 1: Sparse Attention Vectors (SAVs) Overview. We develop a method for extracting features from a generative LMM without finetuning. We first extract a sparse set of attention vectors for each task given a set of few-shot examples, and then, we utilize these attention vectors directly as features for downstream vision-language classification tasks.
Feature extraction is a well-explored field in both vision-only [93, 82, 14] and language-only encoder models [12, 80], but such is not the case with generative models. Most current methods for extracting features from generative models require carefully constructed prompts [40, 102, 66], specialized architectures [50], few-shot prompting [6, 110, 10], and finetuning [63, 7, 109]. Recent work [109] shows that prompting approaches fail to close the gap with encoder VLMs, while finetuning requires training-scale data for every unseen task, which is inefficient. However, generative models still offer the promise of more flexible, truly multimodal features as compared to modality-specific features extracted from CLIP-like models. As such, we are motivated to extract multimodal features from a (frozen) generative LMM, for use in any downstream vision-language classification task.
One source of inspiration for our method is long-standing work in the field of neuroscience that suggests certain areas of the brain are reserved for specific tasks [36, 16] (i.e. functional specificity). Motivated by this idea, we refer to recent interpretability research that has focused on identifying specific heads in transformer-based models that correspond to particular tasks [71]. The most prominent of these methods is a line of work that looks to enhance vision-language capabilities using task vectors [23, 92, 26, 29], which are compact implicit representations of tasks encoded in the activations of a transformer model. While promising, these methods ultimately use these representations to augment a model’s generative capabilities. On the other hand, we seek to use feature representations directly as classifiers. Nevertheless, this intuition from interpretability informs our work on Sparse Attention Vectors (SAVs), which are sparse features in an LMMs activation space that can be directly exploited for few-shot VL classification.
Our method has three steps: First, we extract features (i.e., attention vectors) from the output of each head of the LMM for some few-shot labeled examples ($\approx 20$ per label). Second, we average these attention vectors over the examples in each class and evaluate their accuracy as centroids in a class centroid classifier. We then select the top 20 heads by classification accuracy as our SAVs. In this way, we identify a very sparse set of attention vectors (less than 5% of the total number of heads) that can be used for discriminative tasks. Finally, we perform inference on the given task by doing a majority vote across this sparse set of attention vectors for each new query. This approach requires only few-shot examples at test-time to extract effective multimodal embeddings. An overview is shown in Figure 1.
We summarize our main contributions as follows: (i) We introduce a novel method that yields a sparse set of attention vectors to serve as highly effective features for individual VL classification tasks; (ii) Our method surpasses zero-shot, few-shot, and LoRA fine-tuned baselines across multiple tasks (+7% improvement on average over LoRA on challenging benchmarks like BLINK [17], VLGuard [112], and NaturalBench [43]); (iii) We establish several advantageous properties of our approach, including strong generalization capabilities and favorable scaling characteristics.
Figure 2: Sparse Attention Vectors (SAVs) Detailed View. Our method is broken into the following three parts: (1) Given a set of few shot examples to be classified by a frozen LMM, we extract attention vectors across all heads and layers. (2) These attention vectors are averaged across the set of examples for each class. For each head, we use these mean attention vectors as centroid classifiers, which are then used to select a sparse set of $k$ heads with the highest classification accuracy. (3) Finally, we use these sparse attention vectors to directly classify new inputs via majority vote.
2 Related Works
Controllable Generation for Classification. Controllable text generation in LMMs guides model outputs to meet specific constraints. For classification tasks with generative LMMs, several approaches exist: test-time hard prompting [102, 6] uses prompt engineering or few-shot examples to elicit class label outputs [10, 87, 100, 103, 109]; direct probability analysis of generated class labels [54, 57] for image-text retrieval; and soft prompting methods that finetune learnable tokens [41, 48]. Other techniques include instruction finetuning [101] on labeled data [109, 5] and preference modeling like DPO and RLHF [78, 15, 73]. Our method, however, is finetuning-free and directly selects class labels without preference data. Most related work shows that internal transformer representations called task vectors [91, 22, 26, 29] can encapsulate ICL example tasks. Beyond previous approaches, we use a sparse set of attention vectors directly as VL classification task task features.
Vision-Language Features. Feature extraction seeks useful representations for diverse downstream tasks. Early embedding techniques including autoencoders [4, 81, 38, 61, 62], Word2Vec [65] and GloVe [76] transformed inputs into vector representations, followed by advances in NLP [12, 80, 67, 18] and vision [93, 82, 14]. Recent methods like CLIP and SigCLIP [77, 45, 46, 106, 107, 49] explore multimodal correlations through contrastive learning or sigmoid loss. These representations offer flexibility across tasks [11, 94, 39, 79, 24, 32] and domains [108, 99, 52].
Interestingly, extracting features from generative models poses unique challenges in identifying optimal extraction points. Some approaches finetune encoder VLMs on LLM-generated data [47, 96] or finetune LMMs directly on specific tasks [109, 30]. More efficient methods finetune encoder-decoder representations for better alignment [69, 68, 63, 35]. Finetuning-free approaches use distillation prompts to extract representations from model weights or activations [34, 33, 60, 40], while other methods employ mixture-of-experts [51], expert models [98], or embedding reranking [20].
Thus, current SOTA faces the following challenges: (1) modality-specific rather than multimodal features, (2) requiring finetuning, (3) limited flexibility due to specialized prompts, and (4) dependence on multiple models. Our approach provides effective multimodal embeddings without any gradient-based finetuning and flexibly apply to various VL classification tasks without additional models.
3 Methods
In this section, we outline our approach for using sparse attention vectors from the activation space of a transformer-based large multimodal model (LMM) as features for any VL classification task. The method consists of three main steps: (i) extracting the attention vectors from all attention heads in the model, (ii) identifying a sparse set of vectors based on their ability to consistently return the correct label for some support set of examples, and (iii) using these sparse features to classify new queries. We begin with a formal description of the transformer decoder LLM and its attention mechanism, followed by the detailed methodology for sparse attention vector selection and classification. A detailed view of our method is shown in Figure 2.
3.1 Preliminaries
A transformer-based large language model (LLM) with $L$ layers and $H$ attention heads per layer processes input sequences through multi-head self-attention mechanisms. Each layer combines multiple attention heads to capture different aspects of the input sequence, followed by feed-forward networks for further processing.
Multi-Head Attention. Let $x=\{x_{1},x_{2},\ldots,x_{T}\}$ represent a sequence of input tokens, where $x_{i}$ is the $i^{\text{th}}$ token. For each layer $l\in\{1,\dots,L\}$, the input sequence is projected into queries, keys, and values for each attention head $m\in\{1,\dots,H\}$. Each head performs the following scaled dot-product attention:
$\mathbf{h}_{l}^{m}(x_{i})=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{m}}}\right)V$
where $Q$, $K$, and $V$ are the query, key, and value matrices respectively, and the dimensionality of each head $d_{m}$ which is given by $\frac{d}{H}$ (the embedding dimension divided by the number of heads). We denote $\mathbf{h}_{l}^{m}(x_{i})$ as an attention vector for head $m$ in layer $l$.
The outputs of all heads are concatenated and projected to form the layer output:
$\text{MultiHead}(x_{i})=\text{Concat}(\mathbf{h}_{l}^{1}(x_{i}),...,\mathbf{h}_{l}^{H}(x_{i}))W^{O}$
where $W^{O}$ is the output projection matrix.
In our work, we look to leverage attention vectors for the purpose of vision-language classification tasks. Specifically, the attention vectors are used as latent representations of the inputs to both find attention heads in an LMM suited for a classification task and then perform downstream inference using those selected attention heads. We describe our method in detail in the sections that follow.
3.2 Sparse Attention Vectors
Our key insight is that within the many attention heads and transformer layers of an LMM, there exists a sparse subset that can serve as effective features for vision-language classification tasks. We present a three-step method to identify and utilize these features to build lightweight classifiers.
Step 1: Extracting Attention Vectors. Given a frozen LMM and few-shot examples of sequence-label pairs $\{(x_{1},y_{1}),(x_{2},y_{2}),\dots,(x_{N},y_{N})\}$ we first extract the attention vectors for each sequence $x_{i}$. Specifically, we compute the attention vector $\mathbf{h}_{l}^{m}(x_{i}^{T})$ for head $m$ from layer $l$ for the final token $x_{i}^{T}$. This yields a set of attention vectors $\{\mathbf{h}_{l}^{m}(x_{i}^{T})\mid i=1,\dots,N\}$ for each head $m$ and layer $l$.
Step 2: Identifying Relevant Vectors. The central question is how to identify which attention vectors are naturally suited for the classification task at hand. We evaluate each vector’s ability as a local classifier by computing its performance under a nearest class centroid classifier.
Specifically, for each class $c\in\mathcal{C}$, compute its centroid (or mean) attention vector across the few shot examples:
$\mu_{c}^{l,m}=\frac{1}{|N_{c}|}\sum_{j:y_{j}=c}\mathbf{h}_{l}^{m}(x_{j}^{T})$
where $N_{c}=\{j:y_{j}=c\}$ is the set of indices of examples with label $c$. For each input $x_{i}$, we compute its cosine similarity to each class centroid head:
$s_{l,m}(x_{i},c)=\frac{\mathbf{h}_{l}^{m}(x_{i}^{T})\cdot\mu_{c}^{l,m}}{\|\mathbf{h}_{l}^{m}(x_{i}^{T})\|\|\mu_{c}^{l,m}\|},\quad\forall c\in\mathcal{C}$
Next, we measure the classification accuracy of each head (i.e. local classifier):
$\text{score}(l,m)=\sum_{i=1}^{N}\mathbf{1}[\hat{y}=y_{i}]$
where the nearest class centroid label is given as $\hat{y}=\underset{c\in\mathcal{C}}{\arg\max}\ s_{l,m}(x_{i},c)$, and $\mathbf{1}[\cdot]$ is the indicator function that evaluates to 1 when the condition is true (and 0 otherwise). We denote the set of $k$ top-scoring heads as $\mathcal{H}_{\text{SAV}}$:
$\mathcal{H}_{\text{SAV}}=\{(l,m)\mid\text{score}(l,m)\text{ is among }k\text{ highest scores}\}$
Step 3: Classification with Sparse Attention Vectors.
Given a query sequence $Q$ to classify, we leverage our sparse set of heads $\mathcal{H}_{\text{SAV}}$ for prediction. For each head $(l,m)\in\mathcal{H}_{\text{SAV}}$, we compute the class centroid $\mu_{c}^{l,m}$ closest to the query as follows:
$\hat{y}_{l,m}=\underset{c\in\mathcal{C}}{\arg\max}\ s_{l,m}(Q^{T},c)$
where $s_{l,m}(\cdot,\cdot)$ is defined as in Step 2. Our final class prediction is that of a global classifier that counts the majority vote across all local classifiers (heads) in $\mathcal{H}_{\text{SAV}}$:
$\underset{y\in\mathcal{C}}{\arg\max}\sum_{(l,m)\in\mathcal{H}_{\text{SAV}}}\mathbf{1}[\hat{y}_{l,m}=y]$
This approach reveals a surprising capability of LMMs: with just a few carefully selected attention heads ($|\mathcal{H}_{\text{SAV}}|\ll LH$), we can transform a generative language model into a lightweight vision-language classifier. This finding suggests that classification-relevant features naturally emerge within specific attention heads during model pretraining. In theory, one could replace our local and global classifiers with more complex varients (e.g. linear models, MLPs), something we explore in 3(a).
Image-Text Tasks Image-Only Tasks
Safety VQA I-T Retrieval Classification
Model MHalu VLGuard BLINK VizWiz NB (T) NB (I) NB (G) NB-r (T) NB-r (I) NB-r (G) SC EuroSAT Pets Imagenet-1k Flower CUB
CLIP - - - - - - - 41.8 45.0 23.2 35.3 64.0 88.1 96.1 92.8 97.8
SigLip - - - - - - - 54.5 54.9 31.2 42.7 63.9 98.3 97.6 95.8 98.4
GPT-4o 39.4 45.7 59.0 58.1 64.6 66.4 39.6 65.0 67.0 40.5 48.3 70.5 98.3 98.3 99.0 99.3
LLaVA-1.5 44.3 31.0 38.0 50.0 37.7 43.8 12.7 36.7 42.7 12.2 24.9 26.3 43.0 20.6 51.8 49.0
Instruct-BLIP 44.0 18.5 38.7 34.5 20.2 24.2 4.0 19.5 21.3 1.1 4.7 20.4 56.0 11.0 18.2 33.8
LLaVA-OV-7B 34.7 31.4 45.0 60.4 52.0 53.3 27.0 56.2 58.0 32.1 15.3 66.5 88.1 92.5 83.2 85.3
+4-shot-ICL 25.0 35.0 38.9 47.8 47.6 50.4 22.2 48.2 49.6 31.5 16.1 47.1 63.9 49.0 63.8 60.6
+MTV 37.3 32.9 44.5 61.1 56.2 58.0 30.7 58.1 59.5 33.4 28.7 65.5 88.5 95.9 83.2 85.6
+LoRA 78.3 90.0 47.0 63.1 58.6 60.9 32.4 59.4 60.3 35.4 30.4 85.0 96.8 93.2 91.2 91.8
+SAVs 80.8 94.3 51.8 66.1 60.3 62.3 35.1 72.7 73.0 53.1 37.6 86.7 97.0 97.5 99.6 97.5
+46.1 +62.9 +6.8 +5.7 +8.3 +9.0 +8.1 +16.5 +15.0 +20.5 +22.3 +20.2 +8.9 +5.0 +16.4 +12.2
Qwen2-VL-7B 24.0 26.9 43.3 68.3 53.8 56.6 28.5 60.2 61.9 35.6 24.9 54.7 92.6 84.4 93.7 93.2
+4-shot-ICL 40.4 52.9 37.6 67.1 38.2 41.3 15.2 42.4 45.6 22.7 25.2 29.4 43.5 43.8 59.4 48.4
+MTV 32.3 21.9 41.9 68.5 54.8 57.3 29.7 63.5 64.0 37.0 40.7 52.3 91.7 91.2 94.3 94.1
+LoRA 84.8 87.7 46.3 70.8 55.3 57.4 28.8 65.2 66.1 40.4 38.4 72.9 98.4 96.4 97.1 95.0
+SAVs 85.1 96.0 47.2 68.3 57.6 60.9 32.3 70.0 71.0 42.5 47.5 79.9 98.1 97.6 99.8 98.7
+61.1 +69.1 +3.9 +0 +3.8 +4.3 +3.8 +9.8 +9.1 +13.8 +22.6 +25.2 +5.5 +13.2 +6.1 +5.5
Table 1: Results evaluation on Safety, Visual Question Answering (VQA), Image-Text Retrieval (I-T Retrieval), and Image Classification benchmarks. The best result for each generative model is shown in bold and the second best in underline. We gray out additional baselines. We note that CLIP and SigLIP cannot be evaluating directly on tasks with interleaved image-text queries. $\dagger$ GPT-4o is a close-source model and as such, is shown just as an upperbound (since SAVs are not directly applicable). Key: NB - NaturalBench, SC - SUGARCREPE.
4 Evaluation
We apply SAVs to two state-of-the-art LMMs—LLaVA-OneVision [44] and Qwen2-VL [97]. We also do a rigorous comparison of our method to strong few-shot and finetuning baselines on a variety of different image-text and image-only vision-language classification tasks covering safety, VQA, image-text retrieval, and simple classification.
4.1 Implementation Details
We implemented our approach in PyTorch [75]. We use the official implementations of each model, and all of our experiments can be run on a single NVIDIA A100 GPU. We also use the hyperparameters specified by the original model, which are all tuned for multi-task VL training such as the ones we report in our paper. More details of the implementation is included in the supplementary in Section 8.
Figure 3: Head Visualization. On the top row, we show the top-20 attention head locations for a given task. All the heads are indexed by their (head, layer) position, and the selected heads are highlighted. On the bottom row, we visualize the attention vector of few-shot examples for the top head of the given class with t-SNE clustering [25]. Each point represents the embedding of an input sample.
4.2 Benchmarks
[…]
4.3 Baselines
For our results, we utilized SAVs with 20 examples per label. We compared our method with multiple SOTA baselines, including GPT-4o [72]. As a closed-source model, GPT-4o is presented simply as a strong closed-source baseline. Furthermore, we present the results of classic LMMs, LLaVA-1.5 and InstructBLIP, which have been previously applied to image classification [109]. Zero-shot (ZS) baselines are implemented by querying the model directly and generating a response. For image-text retrieval, ZS uses the SOTA generative scoring method VQAScore [56]. In addition to ZS, we also compare to several test-time and finetuning few-shot methods (all with the exact same sample complexity as SAVs). For instance, we compare to the current SOTA multimodal few-shot method, MTV [29] as well as doing 4-shot ICL and LoRA [28] finetuning for each task.
4.4 Results
Figure 4: Scaling Property of SAVs. Performance of LLaVA-OneVision-7B + SAVs on varying number of few-shot examples per label (left). Performance of of LLaVA-OneVision-7B + SAVs on varying numbers of attention vectors used (right).
Results are shown in Table 1. An advantage of our method is its adaptability to any VL classification task that has image, text, or interleaved image-text inputs. Thus, we apply SAVs to a wide range of tasks in safety, VQA, image-text retrieval and image classification. Furthermore, our approach even significantly improves over SOTA ZS, few-shot, and finetuning methods. One interesting observation is that few-shot ICL consistently deteriorates performance. We hypothesize that instructioning tuning of LMMs disrupts the few-shot prompting capabilities gained during pretraining in favor of structured instructions as shown in [85, 13]. On the other hand, SAVs overcome this by directly operating on the internal activations. Finally, not only is our method successful across a wide-range of tasks, but it also improves on challenging visual perception and compositional reasoning tasks (e.g., BLINK and NaturalBench) that all VL models struggle with. Please refer to our supplementary section in Section 7 for mor results.
4.5 Ablations
We perform a comprehensive ablation study of our method on MHaluBench, NaturalBench, and EuroSAT (see Table 2). For more ablations, please refer to Section 7.1 in the Supp. For all ablations, we use LLaVA-OneVision-7B.
Varying number of examples. In  Figure 4 (left), we examine the impact of varying the number of few-shot examples used in our method. Our primary results in Table 1 indicate that just 20 examples per label are necessary to yield state-of-the-art performance on a variety of VL classification tasks. This ablation shows that accuracy on these tasks can scale with increasing numbers of examples per label.
Varying number of attention vectors. In order to extract attention vectors, we select a very sparse set of heads from hundreds. We vary the number of attention vectors selected in Step 2 of our method and demonstrate that just 20 vectors are enough to realize nearly all of the classification accuracy of our method. Results are shown in Figure 4 (right).
SAVs are flexible to different evaluation strategies. In our method, we leverage class centroid classification as the evaluation method for selecting sparse features. We view this flexibility of the sparsification method to be a key feature of our work. As such, we compare our class centroid classification approach to k-nearest neighbors (KNN) and linear probing. For linear probing, we train a lightweight MLP module for 20 epochs using the top heads’ features. All methods make use of the same 20 examples per label for consistency. Our results in Table 3(a) show that class centroid classification outperforms both KNN classification and is comparable with linear probing.
Comparing heads vs. layers. Based on prior work and transformer-architecture intuition, we treat the attention vectors outputted by the heads as a viable set of features for classification tasks. We verify this intuition by comparing the performance of selecting 2 sparse layers to selecting sparse heads as feature maps for our tasks. As shown in Table 3(b), head-based attention vectors outperform concatenated layer features on all three benchmarks.
4.6 Additional Experiments
(a) Generalization
MHB VLG
Zero-Shot 34.7 31.4
VLG LoRA 34.1 90.0
VLG SAV 55.3 94.3
(b) Interleaved Tasks
MHB NB
CLIP 51.9 1.2
SigLIP 48.6 1.2
SAVs 80.8 35.1
(c) SAVs vs Few-Shot SigLIP
NB ES
SigLIP Few-shot Mean 1.2 88.2
SigLIP Few-shot Cross-Modal 6.8 86.4
SAVs 35.1 86.7
Table 2: Additional SAVs Experiments. We (a) demonstrate the generalization of SAV heads to similar tasks, (b) show the effectiveness of SAV for tasks with interleaved image-text inputs, and (c) compare SAVs with few-shot formulations of CLIP and SigLIP.
(a) Classification Methods
MHB NB ES
Class Centroid 80.8 35.1 86.7
KNN 53.0 11.0 78.1
Linear Probe 82.5 32.9 83.1
(b) Sparse Configurations
MHB NB ES
Sparse Heads 80.8 35.1 86.7
Sparse Layers 79.0 28.4 81.8
Table 3: SAVs Ablations. We perform several ablations to identify the important aspects of our method that contribute to its effectiveness. In particular, we compare the effectiveness of (a) different classification methods and (b) head feature sparsification versus layer feature sparsification. Note: MHB represents MHaluBench, NB represents NaturalBench Group Score, and ES represents EuroSAT. For more ablations, please refer to Section 7.1 in the Supplementary.
In this subsection, we present experiments that demonstrate additional properties and capabilities of SAVs, beyond its use as features for VL classification tasks. Additional experiments can be found in Section 7.2 of the Supplementary. For all experiments, we use LLaVA-OneVision-7B.
Visualizing SAVs. SAVs are both an efficient and interpretable method for leveraging generative LMMs for VL classification tasks. To emphasize this point, we show the selected heads for hallucination detection, relative depth, and image classification in the first row of Figure 3. The visualizations demonstrate both the sparsity and specificity of the SAVs that are used for each task. Unlike prompting and finetuning methods, our approach clearly outlines exactly where in the model’s activation space informative attention vector features are extracted from. We furthermore show that the features extracted from these heads are useful for the given task in the second row of the figure, where we visualize the features outputted by the top selected head for each few-shot sample via t-SNE [25]. The clear clustering of examples of the same label indicates that even with a single head, high-quality features are being selected as SAVs.
Evaluating SAVs generalizability. Here, we ask whether SAVs extracted from one task, can generalize to another similar task. We utilize SAV heads from VLGuard to evaluate on MhaluBench and vice versa. We do a similar approach with LoRA, by applying LoRA finetuned on VLGuard to MHaluBench. Interestingly, our results in Table 2(a) show that SAVs generalize between tasks, while LoRA weights, as expected, overfit to the finetuned task.
Comparing SAVs to CLIP/SigLIP on interleaved image-text tasks. As discussed in Section 1, SAVs are fully multimodal features able to represent inputs that are image-only, text-only, and even interleaved image-text. This is something that is not possible to directly replicate with CLIP and SigLIP models which have separate image and text encoders. Nevertheless, we compare our method to both CLIP and SigLIP on tasks that require interleaved image-text inputs. While SAVs can do this natively, we enable this comparison by concatenating the separate image and text features of CLIP and SigLIP in order to evaluate on MHaluBench and NaturalBench. We find that our method vastly outperforms concatenated CLIP and SigLIP features on both benchmarks as shown in Figure 2(b). This result demonstrates the adaptability of our method to any VL classification regardless of the input’s modality.
Comparing SAVs to few-shot SigCLIP. As SAVs are extracted with few-shot examples, we compare our method to an analogous version of few-shot SigLIP. However, because SigLIP cannot be directly made few-shot, we adapt SigLIP as a few-shot class centroid classifier. One method used is the current SOTA few-shot classification method for encoder models Cross-Modal Adaptation [55]. We apply a second method where we aggregate CLIP/SigLIP embeddings into a mean embedding for each label. Then, just as in SAVs, we perform class centroid classification for each query using our set of mean SigLIP embeddings. We note that for image classification like EuroSAT, only image embeddings are needed, but for MHaluBench, multimodal embeddings are necessary. SAVs are inherently multimodal and so can be flexibly applied to both, but CLIP/SigLIP only have image-only or text-only embeddings. To overcome this, we use SigLIP image features for EuroSAT and concatenate image and text features for MHaluBench. Interestingly, while SiGLIP is comparable to SAVs on EuroSAT, our method vastly outperforms SigLIP in the few-shot setting for MHaluBench, suggesting generalizability of our method for a variety of multimodal tasks that CLIP-like struggle with. Our results are shown in Table 2(c).
5 Conclusion
Our research demonstrates the effectiveness of extracting Sparse Attention Vectors (SAVs) from the heads of an LMM and utilizing them directly for vision-language classification. Our method stands out by using only few-shot examples per label and only less than 1% of the heads to outperform zero-shot, few-shot, and fine-tuned baselines on a variety of image-text and image-only tasks. In addition, SAVs allows generative LMMs to close the gap with closed-source GPT-4o while also being an interpretable method that can generalize to similar tasks. Our ablations reveal the flexibility of using any classification method as a sparsification method for attention vectors and also shows that features are found as outputs of heads rather than layers. Overall, these results show that SAVs is a lightweight, performant, and generalizable method for extending generative LMMs’ multimodal classification abilities. We are encouraged by the outcomes, and anticipate many directions for future work. In addition to methodological improvements, we look forward to the application of SAVs as features for multimodal retrieval, data compression, or more generally as a distilled representation for downstream models.
6 Limitations
Sparse Attention Vectors are a significant step in generalizing the capabilities of generative LMMs to classification tasks. Nevertheless, it is valuable to consider certain limitations of our approach. SAVs are a method that requires access to the model’s internal architecture and so may not be directly applicable to closed-source models like GPT-4 [72] and Gemini [89, 88]. Additionally, some tasks like image-text retrieval [90, 27] can benefit from more fine-grained confidence metrics attached to each label than proportion of voting heads per label. These challenges prompt future work in these directions as well as exciting questions about how to use SAVs as feature embeddings for other tasks.
Acknowledgements.
We would like to thank Abrar Anwar and Tyler Bonnen for helpful feedback and discussions. This project was supported in part by DoD, including PTG and/or LwLL programs, as well as BAIR’s industrial alliance programs.
Supplementary Material
Here we provide additional information about additional experimental results, qualitative examples, implementation details, and datasets. Specifically, Section 7 provides more experiment results, Section 8 provides additional implementation details, and Section 9 provides qualitative visualizations to illustrate our approach.
7 Additional Experiment Results
We begin by presenting several additional ablations (Section 7.1) that further demonstrate the benefits of our SAVs approach. We also present additional results (Section 7.2) on BLINK Splits.
7.1 Additional Ablations
(a) ICL Inputs
MHB NB ES
4-shot 28.3 15.2 29.4
SAVs 82.0 35.1 86.7
(b) Example Robustness
MHB NB ES
MTV 39.6 (2.7) 29.2 (1.2) 65.2 (2.2)
SAVs 83.2 (1.7) 34.8 (.87) 86.4 (1.1)
(c) Noise Robustness
MHB NB ES
2-noisy 82.5 36.1 85.9
5-noisy 81.9 35.6 86.0
10-noisy 50.3 3.3 79.0
Table 4: SAV Additional Ablations. We perform several ablations to identify the important aspects of our method that contribute to its effectiveness. In particular, we evaluate the impact of (a) passing examples in in-context learning format, (b) different examples used, and (c) noisy examples used on the performance of SAVs. Note: MHB represents MHaluBench, NB represents NaturalBench Group Score, and ES represents EuroSAT.
In what follows, we provide additional ablations that further illustrate the benefits of SAVss. For all ablations, we use LLaVA-OneVision-7B.
SAVs using ICL Examples. In our method, we use 20 zero-shot examples as features for discriminative VL tasks. Here, we evaluate the impact of formatting all or some of the examples as few-shot ICL. More concretely, we compare SAVs to (1) a single 20-shot ICL attention vector for each class centroid, and (2) averaging 4 attention vectors of 5-shot ICL examples for each class centroid. Our results, shown in Table 4(a), demonstrate that SAVs are effective for any input format of the examples. However, the best performance is observed when using 20 one-shot examples. This indicates some information is lost when the 20-shots are concatenated into an ICL input while also strengthening the intuition that the attention vectors are good features of individual input examples.
Robustness to examples used. To evaluate the effect of using different sets of examples with our method, we run evaluation using different seeds so that our method sees different examples when extracting SAVs. We compare the performance of SAVs to MTV when running 5 different seeds. We report both the mean and standard deviations of these runs in Table 4(b). We find that MTVs and SAVs are similarly robust to different examples used. This indicates that rather than overfitting to the given examples, SAVs are learning the underlying task.
Robustness to noisy examples. We want to further assess whether SAVs are resilient to noisy examples. We test this by including erroneous examples per class. In other words, for each set of 20 examples per class label, 2, 5, or 10 examples are distractors. We find interestingly that even with 2 or 5 noisy examples, SAVs are still able to achieve comparable performance to SAVs without noise. This result indicates that SAVs are able to average out noise that may be extant in the samples. This property is valuable in cases where it is difficult to ensure correctness of all labeled samples, making SAVs an attractive method for custom tasks with hand-labeled data. Our results from this ablation are shown in Table 4(c).
7.2 Additional Results
(a) Impact of Token Position
MHB NB ES
Last 80.8 35.1 86.7
Middle 49.8 2.4 82.7
First 49.4 0 24.9
(b) Language-Only Tasks
SST-2 MNLI
Zero-shot 88.4 62.7
SAVs 94.5 78.8
(c) Online Learning
MHB NB ES
SAVs 82.0 35.1 86.7
SAVs + O.L. 73.2 29.1 83.8
Table 5: SAV Additional Results. We perform several additional experiments to demonstrate different properties and capabilities of SAVs. In particular, we evaluate the effectiveness of our method (a) when selecting attention vectors from different tokens, (b) on language-only tasks, and (c) when using it in an online learning setting. Note: MHB represents MHaluBench, NB represents NaturalBench Group Score, ES represents EuroSAT, and O.L. represents online learning.
Detailed Split Results. We present detailed results of our method on the BLINK dataset. The results are shown in Table 6.
Model Sim. Cou. Dep. Jig. AS FC SC
LLaVA-OneVision-7B 72.1 22.5 73.4 53.3 52.1 16.9 30.0
LLaVA-OneVision-7B-SAVs 75.0 19.2 78.2 72.0 69.2 43.8 32.1
Qwen2-VL-7B 62.5 23.3 66.1 55.3 47.9 20.0 28.6
Qwen2-VL-7B-SAVs 58.1 26.7 68.5 71.3 57.3 35.4 32.9
Model Spa. Loc. VC MV Ref. For. IQ
LLaVA-OneVision-7B 81.8 51.2 29.7 58.6 32.1 33.3 23.3
LLaVA-OneVision-7B-SAVs 81.8 57.6 31.4 48.9 32.0 54.5 28.7
Qwen2-VL-7B 76.2 49.6 32.0 40.6 42.5 34.1 28.0
Qwen2-VL-7B-SAVs 83.9 56.8 22.7 48.9 32.1 37.9 28.0
Table 6: Detailed Results on BLINK. This table describes the split-level results of our method on all splits of BLINK [17]: Similarity [Sim.], Counting [Cou.], Depth [Dep.], Jigsaw [Jig.], Art Style[AS], Functional Correspondence [FC], Semantic Correspondence [SC], Spatial [Spa.], Localization[Loc.], Visual Correspondence [VC], Multi-View[MV], Reflectance[Rec.], Forensic[For.], IQ-test[IQ]].
Token position selection. Because the last-token of a sequence in a decoder-only LMM attends to all of the prior tokens in an input sequence, it is natural to extract SAVs from the heads of the last token. However, to validate this intuition, we compare the performance SAVs to extract sparse vectors from other tokens (first, middle, and last). Overall, our results in Table 5(a) show that the last token is the best option for selecting heads for SAVs.
SAVs for language-only tasks. While we show the importance of SAVs especially for vision-language tasks, the methodology can be a powerful way to learn tasks in the language-only domain as well. We demonstrate in Table 5(b) the effectiveness of SAVs on two common LLM text classification tasks. The two tasks are SST2[86] as well as MNLI[1]. Excitingly, our results indicate that SAVs can be an effective method of feature extraction to enhance discriminative tasks in the language-only setting as well.
SAVs with online learning. Online learning offers a framework to dynamically adapt predictions based on feedback, but it is traditionally challenging to integrate with deep learning due to the need for updates after each example. However, leveraging the sparse nature of SAVs, we adapt a stochastic online learning method [83] (shown in detail in Algorithm 1) to improve query response accuracy. Specifically, instead of a static majority vote, we employ a randomized weighted voting mechanism that dynamically adjusts weights of individual SAVs based on their correctness over time. This allows the system to prioritize SAVs that consistently perform well given new examples. Our results in Table 5(c) show that SAVs with online learning is not quite performant as our method however. There are a few potential reasons for this. First, our method already optimizes for the quality of the expert voters (i.e. the SAVs). Thus, it is reasonable to consider that additional ordering of these experts is not beneficial. Another simple reason is that online learning methods can be very sensitive and as such different parameters or a slightly different method might be additionally beneficial. Regardless, we encourage future work in this domain.
Algorithm 1 Randomized Weighted Majority Algorithm for SAVs
1:Initialize: Set weights $w_{i}(1)=1$ for all $i\in\{1,\dots,20\}$. Set $\epsilon=\sqrt{\frac{\log d}{T}}$, where $d=20$ is the number of SAVs and $T$ is the total number of queries.
2:for $t=1,\dots,T$ do
3:       Compute selection probabilities $P(i)=\frac{w_{i}(t)}{\sum_{j=1}^{d}w_{j}(t)}$.
4:       Randomly select a SAV $i$ with probability $P(i)$.
5:       Output the prediction of the selected SAV.
6:       Observe the ground truth $y_{t}$.
7:       for each SAV $j\in\{1,\dots,d\}$ do
8:             if SAV $\text{SAV }j\text{ is incorrect}$ then
9:                    Update weight: $w_{j}(t+1)\leftarrow(1-\epsilon)w_{j}(t)$.
10:             else
11:                    $w_{j}(t+1)\leftarrow w_{j}(t)$.
12:             end if
13:       end for
14:       Normalize weights: $w_{j}(t+1)\leftarrow\frac{w_{j}(t+1)}{\sum_{k=1}^{d}w_{k}(t+1)}$.
15:end for
8 Additional Implementation Details
As stated before, we implemented our approach in PyTorch [75] using only the official implememtations and weights of each model. Our implementation precisely follows the steps outlined in Section 3. For the MTV baseline, we follow the method and implementation laid out exactly in the original paper [29]. For our LoRA finetuning baseline, we use the hyperparameters that the respective models (LLaVA-OneVision and Qwen2-VL) used during their instruction finetuning phase. We give more details about the datasets we evaluated on in the following subsections.
8.1 MHaluBench
Dataset. MHaluBench [8] is a dataset that evaluates hallucinations of large multimodal models. Current multimodal models, although they demonstrate remarkable capabilities, have shown hallucinations in a variety of tasks, harming their reliability. MHaluBench evaluates hallucinations by feeding the model with modality-conflicting information. We use the default evaluation method provided in the dataset which is to identify whether this scenario is ”hallucinating” or ”not hallucinating”, and compute the accuracy rate on correctly identified scenarios. We evaluate our model on the image-to-text generation tasks in the dataset, as it is the most common usecase for current multimodal models. The image-to-text generation section of the dataset is focused on Image Captioning and Visual Question Answering tasks, with samples from the MS-COCO 2014 [53] validation set and the TextVQA [84] test set. The generative outputs are compiled from mPLUG [104], LLaVA [59], and MiniGPT-4 [111] to form the core of this dataset.
Inference Details. We use the official source of the code and data. The prompt we use to query the model is ”Is the Claim hallucinating? Answer the question with Yes or No.”
8.2 VLGuard
Dataset. is a vision-language safety instruction-following dataset. This dataset contains four categories of harmful content: Privacy, Risky Behavior, Deception and Hateful Speech. Under these four categories are nine subcategories, which are Personal Data (Privacy); Professional Advice, Political, Sexually Explicit, Violence (Risky Behavior); Disinformation (Deception); Sex, Race and others (Discrimination). The dataset is constructed of images from diverse sources and the instructions are generated by GPT4 [72] with each safe image both safe and unsafe instructions, and each unsafe image a single instruction. The dataset has a train set of 2000 images (977 harmful and 1023 safe) and test set of 1000 images (558 safe and 442 unsafe). VLGuard proposed Attack Success Rate for evaluation, but we reformulate the evaluation process by converting it into a classification problem, which suits our approach better. The reformulation is done by querying the model whether the scenarios is ”harmful” or ”unharmful”. In our implementation we extract the SAVs from the train set and evaluates on the test set. The image-instruction pairs in the dataset can be categorized into three categories: safe image and safe instruction, unsafe image and unsafe instruction, unsafe image and safe instructions, we merged the last two categories into one, as they are harmful in general.
Inference Details. We use the official source of the code and data. The prompt we use to query the model is ”Is the image and instruction harmful? Answer the question with Yes or No.”
Task Query
Jigsaw Which image is the missing part in the first image? Select from the following choices. (A) the second image (B) the third image
Relative Depth Which point is closer to the camera? Select from the following choices. (A) A is closer (B) B is closer
Visual Similarity Which image is most similar to the reference image? Select from the following choices. (A) the second image (B) the third image
Art Style Which image shares the same style as the reference image? Select from the following choices. (A) the second image (B) the third image
Spatial Relation {load question} Select from the following choices. (A) yes (B) no
Multi-View Reasoning The first image is from the beginning of the video and the second image is from the end. Is the camera moving left or right when shooting the video? Select from the following options. (A) left (B) right
Object Localization {load question} Select from the following options. (A) Box A (B) Box B
Forensic Detection Which image is most likely to be a real photograph? Select from the following choices. (A) the first image (B) the second image (C) the third image (D) the fourth image
Visual Correspondence Which point on the second image corresponds to the point in the first image? Select from the following options. (A) Point A (B) Point B (C) Point C (D) Point D
Relative Reflectance Which point has darker surface color, or the colors is about the same? Select from the following choices. (A) A is darker (B) B is darker (C) About the same
Counting How many blue floats are there? Select from the following choices. (A) 0 (B) 3 (C) 2 (D) 1
IQ Test Which one picture follows the same pattern or rule established by the previous pictures? Select from the following choices. (A) picture A (B) picture B (C) picture C (D) picture D
Semantic Correspondence Which point is corresponding to the reference point? Select from the following choices. (A) Point A (B) Point B (C) Point C (D) Point D
Functional Correspondence Which point is corresponding to the reference point? Select from the following choices. (A) Point A (B) Point B (C) Point C (D) Point D
Table 7: Queries for each task in the BLINK dataset.
8.3 BLINK
Dataset. BLINK [17] is a dataset containing multimodal tasks that are intuitive for humans and solvable ”within a blink.” However, these tasks, while straightforward for humans, pose significant challenges for multimodal models. The dataset covers a wide range of visual perception and reasoning abilities, providing a comprehensive evaluation framework. The dataset is formulated as multiple choice questions. We evaluate the models by its accuracy on choosing the right answers for the multiple choice questions. By labeling the choices we essentially convert it into a classfication task.
Among the tasks, Jigsaw tests models’ ability to group and align patterns based on the continuity of color, texture, and shape. Relative Depth evaluates models’ capacity to judge spatial depth between points in an image, while Visual Similarity examines their ability to compare intricate patterns and features. Semantic Correspondence focuses on identifying semantically similar points across images, and Functional Correspondence requires understanding of functional roles in objects. Forensic Detection challenges models to distinguish real images from AI-generated counterparts, emphasizing attention to fine-grained visual details. Multi-View Reasoning, which evaluates spatial understanding by requiring models to deduce camera motion between different viewpoints, and Object Localization, which tests precision in identifying correct bounding boxes in images. Relative Reflectance assesses models’ ability to determine which point has a darker surface color or whether the colors are similar, and Art Style evaluates recognition of stylistic similarities in artworks. Counting measures compositional reasoning in complex scenes with overlapping or occluded objects, and Spatial Relation tests comprehension of relationships like ”left” or ”right.” Finally, the IQ Test assesses pattern recognition and spatial reasoning using visual puzzles, while Visual Correspondence evaluates the ability to identify corresponding points between images.
Inference Details. We use the official source of the BLINK dataset. The prompts we used for different tasks are shown in Table 7.
8.4 NaturalBench
Dataset. NaturalBench [43] is a dataset for Visual Question Answering (VQA). LMMs have shown to be struggling with natural images and queries that can easily be answered by human. NaturalBench is difficult by setting as it require compositionality including to understand complicated relationship between objects and advanced reasoning. The dataset revealed the bias of models preferring the same answers regarding different questions. Each sample from this dataset consists of two questions and images with alternating answers, which prevents the biased models that continuously predicting the same answer regardless of the questions from scoring well. The construction of this dataset is semi-automated as the VQA examples are generated from the previous image-text pairs, which are difficult pairs that cutting edge vision language models failed to match. ChatGPT is used to create questions that have different answers for the two images. We formatted the dataset to give more detailed evaluation. Given that there are two images and two questions (with ””Yes” and ”No” as answer) per example, we divided the results into three sections: ”question accuracy” scoring the model for correctly answering a question for both images, ”image accuracy” scoring the model for correctly answering both questions for an image, and ”group accuracy” scoring the model correctly answering the total four pairs.
Inference Details. We use the official source of the code and data. The prompts we use to query the model are the original questions.
8.5 EuroSAT
Dataset. EuroSAT [21] is a dataset with Sentinel-2 satellite images focusing on the issues of land use and land cover. It is a classification dataset and every image in the dataset is labeled. The dataset covers 10 different classes and 27000 images. The images are diversified as they were taken from all over Europe. It covered 34 countries in Europe, and included images taken all over the years. To improve visibility and clarity, images with low cloud levels are specifically picked. The dataset differed from previous datasets as it covers 13 spectral bands, with visible, near infrared and short wave infrared. The dataset was originally designed for supervised machine learning, but now with the powerful multimodal models we can utilize it as a great tool to test the models’ capabilities to classify, and to discern specific details and intricacies in the images. To better suit the scope of our work, we reformulate the problem into multiple choice questions, with one correct choice and the other 3 randomly selected from the remaining 9 classes.
Inference Details. We use the official source of the data. The prompt we use to query the model is ”What type of remote sensing image does the given image belong to? A. Choice 1 B. Choice 2 C. Choice 3 D. Choice 4”.
8.6 Pets
Dataset. Oxford-IIIT-Pets [74] is a classification dataset consisting 37 different classes of cats and dogs. In the 37 classes, 25 are dogs and 12 are cats, in total there are 7349 images. For each class around 2000 to 2500 images are downloaded from the sources and around 200 are picked, dropping vague examples that are (1) gray scale (2) poorly illuminated (3) having another image portrayed the same animal already (4) animal not centered (5) animal with clothes on it. In our implementation we reformulate the problem into multiple choice questions, with one correct choice and the other 3 randomly selected from the remaining 36 classes.
Inference Details. We use the official source of the data. The prompt we use to query the model is ”What type of animal is in the image? A. Choice 1 B. Choice 2 C. Choice 3 D. Choice 4”.
9 Qualitative Visualizations
We present further qualitative success and failure cases of LLaVA-OneVision-7B-SAVs in Figure 5 and Figure 6.
Figure 5: Additional Example Outputs. Above we show some additional examples on the comparison of our method and zero-shot approach for the MHaluBench, NaturalBench and EuroSAT.
Figure 6: Additional Example Outputs. Above we show some additional examples on the comparison of our method and zero-shot approach for the VLGuard and Oxford Pets.
10 Licenses and Privacy
The license, PII, and consent details of each dataset are in the respective papers. In addition, we wish to emphasize that the datasets we use do not contain any harmful or offensive content, as many other papers in the field also use them. Thus, we do not anticipate a specific negative impact, but, as with any machine learning method, we recommend exercising caution.